Find the specified area either to the left of z=-1.75 or to the right of z=3.00
Answer #1

Z ~ N (0,1)

Required area = either left of Z = -1.75 or to the right of Z =3.00

= P( Z <-1.75 ) + P(Z >3.00)

From normal probability table

P(Z < -1.75 ) = 0.0401 and P(Z > 3.00) = 0.0013

Required area = 0.0401 + 0.0013 = 0.0414

The area either to the left of Z = -1.75 or to the right of 3.00 = 0.0414